User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!

Chair/Presenting: Liz

In Attendance: Ellen, Arwin, Tiege, Lucas, Bryn, Jonny, Ross, Jamie

Apologies: Jesse, Nick, Lucia, Jim




General Issues

  • Swipe Cards need to be activated to turn off/on the alarm at Mt Pleasant.
    • if you're in the email you're fine.
    • probably should check when you're next out there…
  • flogit can't be run twice within an hour
  • Arwin's script is fab for the Yg logs…
  • Potential “troubleshooting” page where the common problems and their fix is listed. One already exists somewhere… We need to rearrange…
  • If you are sleeping make sure people can contact you. Ie probably take the cordless phone into the back room.
  • Do we want a phone numbers page on the wiki?

Issues from Handover notes

* Drives getting stuck at Ke (i.e. r4767)
   * turning it off and on again will most likely fix it... Keep trying. Check it in HMI to see if it is one of the drives is not releasing. ie azimuth/elevation 
   * seems to be a variety of symptoms... Please document this in much detail if it happens to you, including what you tried to fix it with (even if it doesn't work) so we can work out exactly what is the problem!!
* "previous source in the schedule not reached before new source was commanded …" Alarm off? (26m)
    * Fixed. Alarm should sound now.
    * Don't mute/beep these kind of errors because this indicates are really big issue and it needs to be noted. Also, editing the alarm files is a bit iffy...
* What happened with Hb in HOB008?
    * HOB experiments are problematic. So maybe need more notes... Various kinds of errors can occur because things are different.
* excavation work at Ke. fixed?
    * Has access now. New networking/power or something. 
* still having camera problems at Yg for control room
* what happened with eremote control? fixed?
    * Hb needs the updated checklist but is fixed.  
* modules not being loaded. (particularly at Hb) Is it okay to just use whatever is there?
    * Check the VSN for the evening experiment start, matches what is there before 5pm.
    * If you are replacing modules green sticker on middle shelf is okay. DO NOT use the RED sticker ones. Don't do it if there is anything else happening on the mk5.
    * Module allocation page!!
    * Look for youtube video on how to change module by Jim with link on the wiki somewhere.... 
/home/www/auscope/opswiki/data/pages/operations/agenda2016_11_24.txt · Last modified: 2016/11/24 02:32 by Ellen Manning